OpenStudy (anonymous):

Show that cos(4)x =8(cos^(4)x -cos^(2)x)+1

ganeshie8 (ganeshie8):

start by factoring out cos^2 from first two terms

OpenStudy (anonymous):

8cos^2(cos^2x -cosx)+1

ganeshie8 (ganeshie8):

\(\huge 8(cos^4x -cos^2x) + 1\) is this the RHS ?

OpenStudy (anonymous):

yes

OpenStudy (y2o2):

\[\large L.H.S = [ 8\cos^ 2(x)(\cos^2(x) - 1) ]+1 = [ -8\cos^ 2(x)\sin^2(x) ]+1 \] \[\large = [ -2 (2\sin(x)\cos(x))^2 ]+1 = [-2(\sin(2x) ) ^2 ] + 1 \] \[\large = 1-2\sin^2(2x) = \cos^2(2x) - \sin^2(2x) = \cos(4x) = R.H.S\]

ganeshie8 (ganeshie8):

good work @y2o2 :)

OpenStudy (anonymous):

thanks guys, i will look at this later because i am super sleepy

OpenStudy (y2o2):

@ganeshie8 Thank you :)